OpenStudy (anonymous):

solutions to: sqrt(3x)-sqrt(x+1)=sqrt(x-2)

OpenStudy (anonymous):

i can get the solution x=-1 but there is supposed to be an x=3 solution too, but i can't figure out how to get it

OpenStudy (anonymous):

{sqrt(3x)-sqrt(x+1)}^2={sqrt(x-2)}^2 (squaring both sides) 3x+x+1-2(sqrt(3x.(x+1))=x-2 4x+1-2(sqrt(3x.(x+1))=x-2 3x+3=2(sqrt(3x.(x+1)) (3x+3)^2=[2(sqrt(3x.(x+1))]^2 (squaring both sides) 9x^2+9+18x=12x^2+12x 12x^2-9x^2+12x-18x-9=0 3x^2-6x-9=0 x^2-2x-3=0 x^2-(3-1)x-3=0 x^2-3x+1x-3=0 x(x-3)+1(x-3)=0 (x-3)(x+1)=0 x=3 or x=-1

OpenStudy (anonymous):

thank you good sir. i appreciate the help. :)